OpenStudy (anonymous):

PLEASE HELP ME: If ⊗a⊗ =3a-9/2 and ⊗b⊗=6, then what is the value of b?

OpenStudy (anonymous):

is it \[\otimes a\otimes =\frac{3a-9}{2}\]?

OpenStudy (anonymous):

i am going to bet it is then interpret the question as "solve for \(b\), \[\frac{3b-9}{2}=6\]

OpenStudy (anonymous):

Oh okay, thank you!